OpenStudy (anonymous):

4x + 6y = 96 What is the y-intercept? What is the x-intercept?

OpenStudy (anonymous):

i would first change this to y=mx+b to get your y ,and then do the same thing and solve for x to get your x intercept

OpenStudy (anonymous):

These need to be ordered pairs btw.

OpenStudy (callisto):

4x + 6y = 96 =>4x + 6y -96= 0 x-int = -c/a = -(-96/4) = 24 y-int =-c/b = -(-96/6) = 16

OpenStudy (anonymous):

I would start with setting y to 0 to get x 4x+0=96 divide both sides by 96 x=24 Do the same for y

OpenStudy (callisto):

general form : ax+by+c=0 Slope = -a/b x-int = -c/a y-int =-c/b

OpenStudy (anonymous):

If you divide both sides by 96 and simplify, you get x/24 + y/16 = 1. This is the intercept-intercept form. The x-intercept is at (24,0), and the y- intercept is at (0, 16).

OpenStudy (callisto):

wow amazing !!!

OpenStudy (anonymous):

To find x-intercept put y=o you will get x-intercept will be 24 ..........and to find y-intercept put x=0 in this way you will get y-intercept is 16. [4(0)+6y=96 in this way 6y=96 & y=16] you wil get (0,16) [4x+6(0)=96 in this way 4x=96 & x=24] you will get(24,0)

OpenStudy (anonymous):

I like to cut to the chase :^)